I A 35-cm-tall, 5.0-cm-diameter cylindrical beaker is filled to its brim with water. What is the downward force of the water on the bottom of the beaker?
Answer #1

Pressure at the bottom of the beaker is P=P_0+ ho gh

Where P_0 is atmospheric pressure.

Po 101 kPa = 101 * 103 Pa

Density of water ρ-1000 kg/m

Height of cylinder h 35 cm 0.35 m

Radius of cylinder r 2,5 cm 0,025 m

P = Po + Pgh = 101 * 103 + (1000 * 9.81 * 0.35) = 104433.5 Pa

Downward force of water at the bottom of beaker F = PA PTr-) = 104433.5 * π * (0.025)--205N